# AHAMask: Reliable Task Specification for Large Audio Language Models without Instructions

### v1

@@ Conclusion
In this work, we propose AHAMask that simply masks some of the attention heads in an LALM to address the instruction sensitivity problem and achieve reliable task specification. No instruction is further needed once a proper AHAMask is obtained. These masks are efficiently obtained through a lightweight training process. We verify by experiments that applying AHAMask can achieve comparable or even better performance than using general instructions, in diverse scenarios. Our findings also suggest that acoustic “functional pathways” do exist in the attention heads of LALMs, which can be an important inspiration for the interpretability of such models. Future research includes analyzing the partitioning and composability of AHAMask and constructing a general text-to-mask converter for more use cases.
@@ Appendix A Appendix A: Data and Model Details
The training and test data sizes and example outputs for each task can be seen in Table 6, together with detailed information about LALMs. We use official checkpoints for all LALMs. All waveforms are downsampled to 16kHz before feeding the LALMs. For SALMONN, the input waveforms are padded or truncated to 30s, and both its Whisper (Radford et al. 2023) and BEATs (Chen et al. 2023) encoders output high dimensional embedding sequences at 50Hz frame rate. Then, a Q-former (Li et al. 2023) is applied to summarize the acoustic sequences into 88 embeddings for the 30s audio segment before feeding the LLM backbone. For Qwen2Audio models, its encoder compresses 100Hz spectrogram inputs into 25Hz embedding sequences before the LLM backbone. For Qwen2Audio-Instruct, the official chat template is also applied on the audio and text data to enable user-assistant dialogue turns. For SALMONN and Qwen2Audio-Base, the instructions tokens are simply appended after the audio features, such as “<|audio_bos|><|audio|><|audio_eos|> <|instruction|> <|output|>
@@ Appendix C Appendix C: Used Instructions
For reproducibility, please refer to Table 7 for detailed instructions used in the Table 2 and Table 3.
@@ Appendix D Appendix D: Visualization of AHAMasks
Table 6: More dataset and model information for each task. There are two test sets for the ASR task: LibriSpeech (Panayotov et al. 2015) test-clean & test-other. The encoder in Qwen2Audio is initialized on the Whisper-large-v3 model and further tuned, while SALMONN freezes the pretrained Whisper-large-v2 and BEATs-iter3+ models.

### v3

@@ Conclusion
In this work, we propose AHAMask that simply masks some attention heads in an LALM to address the prompt sensitivity problem and achieve reliable task specification without instructions. These masks are efficiently obtained through a lightweight training process. We verify by experiments that applying AHAMask can achieve comparable or even better performance than using general instructions, in diverse scenarios. Our findings also suggest that acoustic “functional pathways” do exist in the attention heads of LALMs, which can be an important inspiration for the interpretability of such models. Future research includes analyzing the partitioning and composability of AHAMask and constructing a general text-to-mask converter for more use cases.
Acknowledgements
This work is supported by Open Project of the Key Laboratory of Artificial Intelligence, Ministry of Education (No. AI202405), and the China NSFC Project (No. 92370206).
@@ Appendix A Appendix A: Data and Model Details
The training and test data sizes and example outputs for each task can be seen in Table 10, together with detailed information about LALMs. We use official checkpoints for all LALMs. All waveforms are downsampled to 16kHz before feeding the LALMs. For SALMONN, the input waveforms are padded or truncated to 30s, and both its Whisper (whisper) and BEATs (beats) encoders output high dimensional embedding sequences at 50Hz frame rate. Then, a Q-former (li2023blip) is applied to summarize the acoustic sequences into 88 embeddings for the 30s audio segment before feeding the LLM backbone. For Qwen2Audio models, its encoder compresses 100Hz spectrogram inputs into 25Hz embedding sequences before the LLM backbone. For Qwen2Audio-Instruct, the official chat template is also applied on the audio and text data to enable user-assistant dialogue turns. For SALMONN and Qwen2Audio-Base, the instructions tokens are simply appended after the audio features, such as “<|audio_bos|><|audio|><|audio_eos|> <|instruction|> <|output|>
@@ Appendix B Appendix B: More Explorations of AHAMask
Task Metric w/ Instr. w/ AHAMask
[…]
Table 5: Performance of AHAMask on composite emotion and gender recognition tasks.
Task Metric w/ Instr. w/ AHAMask
[…]
Table 6: Evaluation of single task performance on SALMONN under composite scenarios and LLM-corrected formats. The setup is similar to Table 3, which does not employ an LLM to repair the LALM outputs. We omit the IFR metric here because it is 100% after LLM format correction. WER and ACC are for ASR and GR tasks, respectively. “Instr.” means “instructions”.
In this section, we show more experimental results of AHAMask on SALMONN.
Additional Composite Tasks
We also evaluate AHAMask on other composite tasks besides ASR and GR. Table 5 shows the results on SER and GR tasks with IEMOCAP training and test sets. These results are similar to Table 3, which provide stronger evidence on the validity of AHAMask.
Performance of Single Tasks under Composite Scenarios
In Table 3, we evaluated LALMs on composite tasks, either with instructions or AHAMask. In that comparison, we calculated instruction following rates, and the metrics on single tasks only within instruction-following samples. Here, we consider using an LLM to reconstruct the required format for those instruction non-following samples, and measure the performance of single tasks on all the samples. This LLM is instructed to read the LALM outputs that do not follow the required format, infer the outputs for each single task, and assemble them into the required format. When a task is missing in the LALM ouput, the LLM is asked to use an empty string for that task. We use OpenAI gpt-4o-mini API which exhibits 100% instruction following rate on the tested samples. Then, we obtain the task-specific metrics in Table 6.
These results show that the advantage of AHAMask still remains with LLM format corrections. Compared to Table 3, all metrics become worse, because certain tasks are usually missing when the format is not satisfied. For the GR+“|”+ASR composite task, the ASR WER of AHAMask is slightly higher than that of official SALMONN with instructions. We identify that it is because of more deletion errors, i.e. the ASR task is usually missing among the 0.88% instruction non-following samples from AHAMask.
On the Generalizability of AHAMask
Test Set w/ Instruction w/ AHAMask
LibriSpeech (test-clean)
96.79 98.05
IEMOCAP (Session 5)
88.64 91.62
TEDLIUM (test)
93.83 95.65
CommonVoice11 (test)
50.73 89.96
VoxCeleb1 (test)
95.90 98.75
Table 7: Out-of-domain evaluation on GR task on SALMONN. Metric is ACC (%). Parentheses denote dataset partition. Note that we use the same AHAMask trained on LibriSpeech train-clean-100 in the last column.
In almost all experiments in this paper, each AHAMask is obtained by training only on one dataset. Ideally, since each AHAMask is supposed to activate a function of the LALM, it should generalize well across different domains. However, as it is acquired by data-driven neural network training in practice, its generalizability also needs verification.
We first observe that for SALMONN, on the GR task, AHAMask exhibits good generalizability on out-of-domain data, despite only being trained on LibriSpeech. Results in Table 7 demonstrates that the AHAMask for GR task on SALMONN performs well on IEMOCAP, TEDLIUM (rousseau2012tedlium), CommonVoice11, and VoxCeleb1 test sets.
Then, on ASR task, we find the generalizability of AHAMask still has room of improvement. When evaluating on the out-of-domain TEDLIUM test set, we observe noticeable discrepancy between AHAMask (trained on LibriSpeech only) and the official instructed model. When we use mixed training set for AHAMask, such as a balanced mix from LibriSpeech and CommonVoice11, we obtain WER scores that match well with the instructed model. Those results are shown in Table 8. This may indicate that AHAMask can still be affected by the training data domains. When trained on a single domain, it may capture a more fine-grained functionality (such as transcribing clean audiobook recording). Therefore, using more diverse data domains, like the way original LALMs are trained, is always beneficial. It remains a future work on evaluating and improving the generalizability of AHAMask, so that it can manifest the general functionality of LALMs more steadily.
AHAMask Training Data w/ AHAMask w/ Instr.
LibriSpeech (all, 280k) 4.01 3.65
LibriSpeech (50k)
+CommonVoice11 (50k)
3.67
Table 8: Out-of-domain evaluation on ASR task on SALMONN. Metric is WER (%). Parentheses denote the number of training samples. The test set here is TEDLIUM (test), which is unseen for both the original SALMONN model and AHAMask. “Instr.” means “instructions”.
@@ Appendix C Appendix C: Used Instructions
For reproducibility, please refer to Table 11 for detailed instructions used in the main text. Please also refer to the supplementary files for detailed instructions we use for SALMONN in the instruction sensitivity experiment.
@@ Appendix D Appendix D: Visualization of AHAMasks
Table 10: More dataset and model information for each task. There are two test sets for the ASR task: LibriSpeech (librispeech) test-clean & test-other. The encoder in Qwen2Audio is initialized on the Whisper-large-v3 model and further tuned, while SALMONN freezes the pretrained Whisper-large-v2 and BEATs-iter3+ models.